9. A study group is to be selected from a group of 16 students. The group consists of 5 freshmen, 7 sophomores, and 4 juniors.

   a) In how many ways can a group of 6 students be selected from this group?

   b) If the study group is to consist of 2 freshmen, 3 sophomores, and 1 junior, how many different groups are possible?

   c) What is the probability that a group of 6 students will consist of 2 freshmen, 3 sophomores and 1 junior?
Answer #1

a)
total 16 students
6 can be selected in 16C6 = 8008 ways

b)
2 freshmen can be selected in 5C2 ways
3 sophomores can be selected in 7C3 ways
1 junior in 4C1 ways

Hence required combinations are 5C2 * 7C3 * 4C1 = 1400

c)
requried probability = 1400/8008 = 0.1748